OpenStudy (cookie_2046):

The information given represents lengths of sides of a right triangle with c the hypotenuse. Find the correct missing length to the nearest hundredth. A calculator may be helpful. a = 18, b = 26, c = ? A. 31.16 B. 31.46 C. 31.62 D. 31.78

OpenStudy (anonymous):

\[a^2+b^2=c^2 \] so \[18^2+26^2=c^2\]